# Multigraded Stillman’s Conjecture

### v1

@@ 1. Introduction
The following example demonstrates that allowing the grading group to have infinite descending chains (i.e. the divisibility order is not well-founded) permits counterexamples to a multigraded Stillman bound on projective dimension.
@@
such that for all variables $x\in X$, $x\in S_{g}$ for some $g\in\Gamma$. If $s\in S_{g}$ we say that $s$ is homogeneous of degree $g$ and write $\deg(s)=g$. We say that $S$ is connected if $S_{0}=k$. The (grading) support of $S$ is the submonoid of $\Gamma$ generated by the degrees of all the monomials along with the identity. We say a (possibly inhomogeneous) ideal has degree sequence bounded by $\bm{d}=(d_{1},\dots,d_{n})\in\Gamma^{n}$ if the ideal can be generated by elements $f_{1},\dots,f_{n}$ whose monomials have degree less than $d_{1},\dots,d_{n}$. Any submonoid $\Lambda\subseteq\Gamma$ has bounded factorization if it is impossible to express an element in $\Lambda$ as an arbitrarily large sum of other elements in $\Lambda$. Our main theorem shows that for a fixed $\Lambda\subseteq\Gamma$, any family of connected $\Gamma$-graded polynomial rings with support contained in $\Lambda$ has Stillman bounded projective dimension if and only if $\Lambda$ has bounded factorization.
@@
After fixing any particular polynomial ring $S$ and ideal $I$, the projective dimension of $S/I$ does not change as one varies the grading. Hilbert’s syzygy theorem guarantees that the projective dimension of every ideal in $S$ is bounded by the (hopefully finite) number of variables in $S$. The classical Stillman’s conjecture ensures a bound shared across the entire family of standard graded polynomial rings at the cost of restricting the desired class of ideals $I$ to those with a particular degree sequence. Theorem 1.2 asserts that if you also allow your grading to vary, under mild conditions you must only further fix $\Lambda$ for Stillman uniformity. Example 1.1 fails to have Stillman bounded projective dimension exactly because the supports of the family fail to lie in a bounded factorization monoid. Letting $\deg(x)=\deg(y)=0$ and $\deg(z_{i})=d$ in the same example shows that non-connected gradings do not have a Stillman bound. We emphasize the fact that the bound in Theorem 1.2 works over families of different gradings with the following example.
Example 1.3.
After finding the right condition on $\Lambda$, the proof of Theorem 1.2 ends up being very short. The key intuition is to leverage the known Stillman bound for the ${\mathbb{Z}}_{+}$-graded case by using bounded factorization to guarantee a nice regrading by ${\mathbb{Z}}_{+}$ for the “if” direction, and to construct an explicit counterexample using only non-bounded factorization for the “only if” direction. Even in the $\Lambda={\mathbb{Z}}_{+}$ case proven by Ananyan and Hochster [AnanyanHochster2016], Theorem 1.2 says something new; the same Stillman bound holds for families of varying ${\mathbb{Z}}_{+}$-gradings, with the bounds potentially getting tighter for polynomial rings whose support refines ${\mathbb{Z}}_{+}$. That is, by equipping a ring with a finer-graded structure, we can get better bounds for a given ideal, as shown in the following example.
Example 1.4.
Under this new grading, the degree sequence of $\langle f,g,h\rangle$ is $\bm{e}=\{(3,0,0),(0,3,0),(0,0,3)\}$. This grading is finer in the sense that the family of ideals that are degree $\bm{e}$ is strictly contained in the larger family of ideals that are degree $\varphi(\bm{e})=\bm{d}$, where $\varphi$ is the natural flattening map summing the entries. Therefore one would expect that $N(\bm{e})<N(\bm{d})$ and in this case, this is true; the Stillman bound under this grading is exactly the projective dimension of $3$.
In Example 1.4, the ${\mathbb{Z}}^{3}$-grading reveals that the variables showing up among $f$, $g$, and $h$ are disjoint and therefore $f,g,h$ form a regular sequence. In general, if the bound in Theorem 1.2 were made effective (e.g. as in the standard graded case [kazhdan2020properties]), one could find better Stillman bounds for a fixed polynomial ring $S$ and ideal $I$ by varying the grading. That is, given a fixed ${\mathbb{Z}}_{+}$-graded $S$ and ideal with degree sequence $\bm{d}$, you may consider equipping $S$ with various positive $\Gamma$-gradings with flattening maps $\varphi:\Lambda\to{\mathbb{Z}}_{+}$ to obtain the bound
[…]
which Example 1.4 shows can be tight.
Open problems
Here are some open problems raised by our work:
  • When is the bound in Equation (1) tight?
Acknowledgements
@@ 2. $\Gamma$-graded Commutative Algebra
$g\leq_{\Lambda}h\iff g+q=h$ for some $q\in\Lambda$.
@@ 3. Main Results and Corollaries
For the forward direction, suppose that $\Lambda$ does not have bounded factorization. This means there exists some $d\in\Lambda$ such that for every $b\in{\mathbb{N}}$ we can find another number $B$ larger than $b$ that is the length of a factorization $d=d_{1}+\dots+d_{B}$, where $d_{1},\dots,d_{B}$ are nonzero degrees in $\Lambda$. Grouping terms arbitrarily, we can take $B=b$. For every $b$ we can define a polynomial ring $S_{b}=k[x_{1},\dots,x_{b},y_{1},\dots,y_{b}]$ with $\Lambda$-grading defined by setting $\deg(x_{i})=\deg(y_{i})=d_{i}$. Then with $f_{1},f_{2},f_{3}$ defined as in Example 3.2, we have that $f_{1}$ and $f_{2}$ are homogeneous of degree $d$, while $f_{3}$ is inhomogeneous but with the degree of each monomial bounded by $2d$. Letting $I_{b}=\langle f_{1},f_{2},f_{3}\rangle$, Example 3.2 confirms that $\operatorname{pdim}_{S_{b}}(S_{b}/I_{b})=b+2$ can grow arbitrarily big, so there cannot be Stillman bounded.
@@ 4. Applications and Examples
As discussed in the introduction, some $\Gamma$-gradings under consideration are exotic enough to fail to admit a flattening map $\Lambda\to{\mathbb{Z}}_{+}$. The following is one such example with Stillman bounded projective dimension:
@@
Let $S=k[x_{0},x_{1},x_{2},\ldots]$ and define $p_{n}$, where $n\in\mathbb{N}$, to be the $n$th prime natural number. Define the degree of any constant to be 00, $\deg(x_{0})=1\in\mathbb{Q}$, and $\deg(x_{n})=n+\frac{1}{p_{n}}\in\mathbb{Q}$. Let $\Lambda\subseteq\mathbb{Q}$ be the $\mathbb{Q}$-support of this grading and then let $\Gamma$ be the subgroup of $\mathbb{Q}$ spanned by $\Lambda$. Then the grading on $S$ is connected and thus $\Lambda$ is pointed. Given any $q\in{\mathbb{Q}}$ it is clear that $\{n\mid n+\frac{1}{p_{n}}\leq q\}$ is finite, so $\Lambda$ has bounded factorization. Therefore due to Theorem 1.2, $S$ has Stillman bounded projective dimension. Interestingly, there is no flattening homomorphism $\Lambda\to\mathbb{Z}_{+}$. Indeed if $\varphi$ is such a map, since $1$ and $\frac{1}{p_{n}}$ are contained in $\Gamma$ for all $n$, we have that $\varphi(1)=p_{n}\varphi(\frac{1}{p_{n}})$ for all $n$, which is impossible as $\varphi$ takes values in the integers.

### v2

# Semigroup Graded Stillman’s Conjecture

@@ 1. Introduction
The following example demonstrates that allowing the grading group to have infinite descending chains (i.e. the divisibility order is not well-founded) can prevent Stillman bounds on projective dimension.
@@
such that for all variables $x\in X$, $x\in S_{g}$ for some $g\in\Gamma$. If $s\in S_{g}$ we say that $s$ is homogeneous of degree $g$ and write $\deg(s)=g$. We say that $S$ is connected if $S_{0}=k$. The (grading) support of $S$ is the submonoid of $\Gamma$ generated by the degrees of all the monomials along with the identity. If $S$ is connected, this submonoid inherits a partial order to be described in (1). We say a (possibly inhomogeneous) ideal has degree sequence bounded by $\bm{d}=(d_{1},\dots,d_{n})\in\Gamma^{n}$ if the ideal can be generated by elements $f_{1},\dots,f_{n}$ whose monomials have degree less than $d_{1},\dots,d_{n}$. Any submonoid $\Lambda$ of the grading group $\Gamma$ has bounded factorization if it is impossible to express an element in $\Lambda$ as an arbitrarily large sum of other elements in $\Lambda$. Our main theorem shows that for a fixed $\Lambda\subseteq\Gamma$, any family of connected $\Gamma$-graded polynomial rings with support contained in $\Lambda$ has Stillman bounded projective dimension if and only if $\Lambda$ has bounded factorization.
@@
After fixing any particular polynomial ring $S$ and ideal $I$, the projective dimension of $S/I$ does not change as one varies the grading. Hilbert’s syzygy theorem guarantees that the projective dimension of every ideal in $S$ is bounded by the (hopefully finite) number of variables in $S$. The classical Stillman’s conjecture ensures a bound shared across the entire family of standard graded polynomial rings at the cost of restricting the desired class of ideals $I$ to those with a particular degree sequence. Theorem 1.2 asserts that if you also allow your grading to vary, under mild conditions you must only further fix $\Lambda$ for Stillman uniformity. Example 1.1 fails to have Stillman bounded projective dimension exactly because the supports of the family fail to lie in a bounded factorization monoid. Letting $\deg(x)=\deg(y)=0$ and $\deg(z_{i})=d$ in the same example shows that non-connected gradings also fail to have a Stillman bound.
After finding the right condition on $\Lambda$, the proof of Theorem 1.2 ends up being very short. The key intuition is to leverage the known Stillman bound for the ${\mathbb{Z}}_{+}$-graded case by using bounded factorization to guarantee a nice regrading by ${\mathbb{Z}}_{+}$ for the “if” direction, and to construct an explicit counterexample using only non-bounded factorization for the “only if” direction. Even in the $\Lambda={\mathbb{Z}}_{+}$ case proven by Ananyan and Hochster [AnanyanHochster2016], Theorem 1.2 says something new; the same Stillman bound holds for families of varying ${\mathbb{Z}}_{+}$-gradings, with the bounds potentially getting tighter for polynomial rings whose support refines ${\mathbb{Z}}_{+}$. We end by discussing some other applications and further examples in Section 4.
@@ 2. $\Gamma$-graded Commutative Algebra
(1) $g\leq_{\Lambda}h\iff g+q=h\text{ for some }q\in\Lambda.$
@@ 3. Main Results and Corollaries
For the forward direction, suppose that $\Lambda$ does not have bounded factorization. This means there exists some $d\in\Lambda$ such that for every $b\in{\mathbb{N}}$ we can find another number $B$ larger than $b$ that is the length of a factorization $d=d_{1}+\dots+d_{B}$, where $d_{1},\dots,d_{B}$ are nonzero degrees in $\Lambda$. Grouping terms arbitrarily, we can take $B=b$. For every $b$ we can define a polynomial ring $S_{b}=k[x_{1},\dots,x_{b},y_{1},\dots,y_{b}]$ with $\Lambda$-grading defined by setting $\deg(x_{i})=\deg(y_{i})=d_{i}$. Then with $f_{1},f_{2},f_{3}$ defined as in Example 3.2, we have that $f_{1}$ and $f_{2}$ are homogeneous of degree $d$, while $f_{3}$ is inhomogeneous but with the degree of each monomial bounded by $2d$. Letting $I_{b}=\langle f_{1},f_{2},f_{3}\rangle$, Example 3.2 confirms that $\operatorname{pdim}_{S_{b}}(S_{b}/I_{b})=b+2$ can grow arbitrarily big, so there cannot be a Stillman bound.
@@ 4. Applications and Examples
We emphasize that the bound in Theorem 1.2 works over families of different gradings with the following example.
Example 4.1.
One possible application of Theorem 1.2 is that by equipping a ring with a finer-graded structure, we can get better Stillman bounds for a given ideal. We demonstrate this with the following example.
Example 4.2.
Under this new grading, the degree sequence of $\langle f,g,h\rangle$ is $\bm{e}=\{(3,0,0),(0,3,0),(0,0,3)\}$. This grading is finer in the sense that the family of ideals that are degree $\bm{e}$ is strictly contained in the larger family of ideals that are degree $\varphi(\bm{e})=\bm{d}$, where $\varphi$ is the natural additive map summing the entries. Therefore one would expect that $N(\bm{e})<N(\bm{d})$ and in this case, this is true; the Stillman bound under this grading is exactly the projective dimension of $3$. This is because if $\langle f,g,h\rangle$ has degree sequence $\bm{e}$, the variables showing up among $f$, $g$, and $h$ are pairwise disjoint and therefore $f,g,h$ forms a regular sequence.
In general, if the bound in Theorem 1.2 were made effective (e.g. as in the standard graded case [kazhdan2020properties]), one could find better Stillman bounds for a fixed polynomial ring $S$ and ideal $I$ by varying the grading. That is, given a fixed ${\mathbb{Z}}_{+}$-graded $S$ and ideal with degree sequence $\bm{d}$, you may consider equipping $S$ with various positive $\Gamma$-gradings with homomorphisms $\varphi:\Lambda\to{\mathbb{Z}}_{+}$ to obtain the bound
[…]
which Example 4.2 shows can be tight.
@@
As discussed in the introduction, some $\Gamma$-gradings under consideration are exotic enough to fail to admit a homomorphism $\Lambda\to{\mathbb{Z}}_{+}$. The following is one such example with Stillman bounded projective dimension:
@@
Let $S=k[x_{0},x_{1},x_{2},\ldots]$ and define $p_{n}$, where $n\in\mathbb{N}$, to be the $n$th prime natural number. Define the degree of any constant to be 00, $\deg(x_{0})=1\in\mathbb{Q}$, and $\deg(x_{n})=n+\frac{1}{p_{n}}\in\mathbb{Q}$. Let $\Lambda\subseteq\mathbb{Q}$ be the $\mathbb{Q}$-support of this grading and then let $\Gamma$ be the subgroup of $\mathbb{Q}$ spanned by $\Lambda$. Then the grading on $S$ is connected and thus $\Lambda$ is pointed. Given any $q\in{\mathbb{Q}}$ it is clear that $\{n\mid n+\frac{1}{p_{n}}\leq q\}$ is finite, so $\Lambda$ has bounded factorization. Therefore due to Theorem 1.2, $S$ has Stillman bounded projective dimension. Interestingly, there is no homomorphism $\Lambda\to\mathbb{Z}_{+}$. Indeed if $\varphi$ is such a map, since $1$ and $\frac{1}{p_{n}}$ are contained in $\Gamma$ for all $n$, we have that $\varphi(1)=p_{n}\varphi(\frac{1}{p_{n}})$ for all $n$, which is impossible as $\varphi$ takes values in the integers.
Open problems
Here are some open problems raised by our work:
  • When is the bound in Equation (2) tight?
Acknowledgements